OpenStudy (anonymous):

y=-x-2 , y=-2x^2+x-3 solve algebraically please help!!!

OpenStudy (anonymous):



OpenStudy (anonymous):

solve \[-x-2 =-2x^2+x-3 \]

OpenStudy (anonymous):

i got -2x^2+2x-1=0 but i dont know where to from there

OpenStudy (anonymous):

write is all on one side of the equal sign as \[2x^2-2x+1=0\]

OpenStudy (anonymous):

or what you had, but it is usually easiest to work with a positive leading coefficient they are the same equation though so it doesn't really matter if you are luck it factors if not, you have to use the quadratic formula

OpenStudy (anonymous):

i dont think it can factor so i would i use quadratic formula?

OpenStudy (anonymous):

actually there are no real number solutions to this equation

OpenStudy (anonymous):

when you attempt the quadratic formula \(b^2-4ac=4-8=-4\) so the solutions are complex numbers